OpenStudy (anonymous):

what is the ratio of the perimeter of a semicircular region to the perimeter of a circular region with the same radius?

OpenStudy (anonymous):

I want to say that it is 1:2 (or 1/2).

OpenStudy (anonymous):

the answer is 2+pi to 2 pi i just dont know how to get to it

OpenStudy (anonymous):

The only reason it's not 1:2 is because the semicircle also has a diameter included. See attachment. Call the radius of each "r." First, the semicircle: The length of the diameter is 2r. You need to add this to half the circumference. The circumference of a circle is 2(pi)r, so half the circumference is pi(r). This means that the perimeter of the semicircle is \[2r+\pi r\]Factor out a pi to get\[C_1=r(2+\pi)\] Next, the circle: The circumference is all we need here. \[2\pi r\]Or alternatively, \[r(2\pi)\] Divide C1 by C2 to get your ratio. \[r(2+\pi)/[r(2\pi)]\]Simplify to get your final ratio of \[(2+\pi)/(2\pi)\]

OpenStudy (anonymous):

Thanks, jabberwock; I didn't think to take the semicircle's diameter into consideration.

OpenStudy (anonymous):

I thought that, too at first :(. Then she wrote the answer and I had to back-solve it.

OpenStudy (anonymous):

Good work!